# NeuSemSlice: Towards Effective DNN Model Maintenance via Neuron-level Semantic Slicing

### v1

@@ Abstract.
Deep Neural networks (DNNs), extensively applied across diverse disciplines, are characterized by their integrated and monolithic architectures, setting them apart from conventional software systems. This architectural difference introduces particular challenges to maintenance tasks, such as model restructuring (e.g., model compression), re-adaptation (e.g., fitting new samples), and incremental development (e.g., continual knowledge accumulation). Prior research addresses these challenges by identifying task-critical neuron layers, and dividing neural networks into semantically-similar sequential modules. However, such layer-level approaches fail to precisely identify and manipulate neuron-level semantic components, restricting their applicability to finer-grained model maintenance tasks.
@@
For semantic-aware model maintenance tasks, we provide a series of novel strategies based on semantic slicing to enhance NeuSemSlice. They include semantic components (i.e., critical neurons) preservation for model restructuring, critical neuron tuning for model re-adaptation, and non-critical neuron training for model incremental development. A thorough evaluation has demonstrated that NeuSemSlice significantly outperforms baselines in all three tasks.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8
@@ 1. Introduction
In the evolving landscape of artificial intelligence, DNNs represent a dynamic shift towards data-driven programming paradigms, extensively adopted across various fields (Kollias et al., 2018; Bansal et al., 2018; Théate and Ernst, 2021). Analogous to how traditional software evolves to meet shifting requirements and enhance functionality, effective maintenance of DNNs is crucial for combating model drift and ensuring sustained accuracy and reliability as they are continuously exposed to new and changing data. Unlike traditional software systems characterized by their modular and discrete structures, DNNs are typically designed as cohesive, monolithic entities. The unique architecture introduces distinct and challenging maintenance tasks, including model restructuring, re-adaptation, and incremental development. They are pivotal elements for the operational robustness and longevity of neural network-based systems in a rapidly advancing technological milieu (AI model maintenance: ensuring accuracy, scalability, and reliability, 2023; AI & ML Model Maintenance, 2023; AI Maintenance: A Robustness Perspective, 2023).
@@
Despite various proposals in the literature for maintaining DNN models like the DeepArc framework (Ren et al., 2023), these methods continue to face several limitations. One major challenge with traditional strategies, which depend largely on periodical model retraining, is their ineffectiveness in combating the unavoidable problem of model drift. DeepArc addresses this problem by modularizing neural networks, organizing layers into modules based on their semantic similarities. This strategy streamlines maintenance tasks, facilitating the pruning and fine-tuning of specific modules or layers for more efficient updates. However, this approach is constrained by its focus on layer-level manipulation. A more refined strategy that decomposes semantic components at a finer level, such as individual neurons, could enhance model restructuring and adaptation. Specifically, selectively pruning and tuning a small number of neurons might offer better effectiveness. Furthermore, the incremental development of models (Wang et al., 2024) requires preserving the semantics and functions of previous versions while introducing new features. Given that semantics and functions in DNNs are primarily structured around neurons (Du et al., 2019; Xie et al., 2022; Zhang et al., 2020b; Li et al., 2021; Shrikumar et al., 2017), the method of identifying and updating whole layers falls short of preserving the original semantics or functions during the incremental development process.
@@
Acknowledging the existing limitations, our goal is to develop techniques that enhance DNN model maintenance. Drawing inspiration from the significant benefits of dynamic slicing in the maintenance and evolution of traditional software (Beszédes et al., 2001; Jiang et al., 2014), we adopt similar slicing techniques for DNNs (Zhang et al., 2020a; Li et al., 2023; Xie et al., 2022; Sun et al., 2022; Liu et al., 2018). Specifically, dynamic slicing excels in identifying and isolating essential code segments for implementing changes without causing unintended effects. By analogy, DNN slicing provides a systematic way to identify and separate semantic components at the neuron level. This approach is poised to greatly enhance DNN model maintenance, allowing for precise modifications and improvements with a focus on finer detail.
@@
The semantic slicing module is designed to accurately identify critical neurons across the entire search space by employing a category-specific [^1] To be precise, we identify critical neurons within the network model that have a significant impact on the output of each output neuron, where different categories refer to different output neurons. For example, in classification tasks, categories can be determined based on the labels in the dataset; in regression tasks, it can be considered a single-category task; in other types of tasks, different output neurons are treated as different categories. strategy for optimal performance. Specifically, we compute neuron contribution scores (Pei et al., 2017; Ish-Horowicz et al., 2019; Xie et al., 2022; Li et al., 2023; Molchanov et al., 2019) for each neuron category individually, selecting a specific range of critical neurons for each category before combining them across the entire dataset. Given the potential variability in the optimal range of critical neurons across different categories and layers—which may hinder effective combination—we refine this process to focus on neurons that encapsulate the layer’s full semantic range. This is achieved by assessing the similarity between the characteristics of these chosen neurons and the entire layer. The subsequent merging strategy further enhances the utility and relevance of the identified critical neurons.
The semantic-aware model maintenance module, utilizing semantic slicing, provides the following maintenance tasks for DNN models: ❶ Model Restructuring, where only critical neurons are retained to balance predictive precision with reduced storage needs. ❷ Model Re-adaptation, focusing on training critical neurons to address prediction errors, thus simplifying the process by tuning fewer parameters. ❸ Model Incremental Development, which stabilizes critical neuron parameters to prevent catastrophic forgetting during the integration of new knowledge, facilitating incremental development without extensive reliance on the original dataset.
@@
We conduct extensive experiments to evaluate the performance of NeuSemSlice. For model restructuring, NeuSemSlice enhances the efficiency of existing model compression techniques by an average of 13.46%, while also improving compression rate. For model re-adaptation, NeuSemSlice only trains an average of 61.49% of the parameters, enhancing both accuracy and efficiency. For incremental model development, NeuSemSlice surpasses all other methods, boosting average accuracy by over 40% compared to baseline.
@@
Contributions. The contribution of this paper is summarized as follows: ❶ We propose NeuSemSlice, including semantic slicing for finer-grained neuron-level decomposition, and semantic-aware model maintenance, offering novel and effective strategies for semantic-aware DNN model maintenance tasks. ❷ We detail a sophisticated semantic slicing technique that employs a category-specific strategy to accurately identify and combine critical neurons across the dataset, refining the process of neuron selection based on semantic range to optimize model performance. ❸ We tackle the challenging maintenance tasks via the semantic-aware model maintenance module, which significantly advances model restructuring, adaptation, and incremental development tasks. ❹ We conduct extensive experiments on three model maintenance tasks, showcasing the superior performance of NeuSemSlice.
@@ 2.1. Deep Neural Networks
In this work, we mainly focus on the classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X}$ is a set of inputs and $\mathcal{Y}$ is a set of classes. Given an input $x\in\mathcal{X}$, we use $f^{l}(x)$ to represent the internal features extracted by the layer $l$ and $f_{n}(x)$ to represent the internal features extracted by the neuron $n$ (i.e., the output values of neurons at $l$). $\hat{y}$ represents the DNNs output (i.e., $f(x)$) and $y$ is the label of the given input $x$.
@@ 2.2. DNN Model Maintenance
This concept, akin to the established practices of traditional software maintenance, addresses the evolving needs within the software engineering community for more efficient and adaptable DNN systems. DeepArc (Ren et al., 2023) has introduced a new neural network architecture strategy to cut model maintenance costs. This approach identifies semantically similar layers as redundant, allowing for model restructuring by pruning these layers. Only non-redundant layers are fine-tuned on the target data, based on calculated semantic similarities of layer representations: Given a dataset $D$, the matrix $M^{l_{i}}$ is denoted as the similarity matrix for the $i$-th network layer $l_{i}$, where
@@
with $x_{a}$ and $x_{b}$ representing any two inputs from the dataset $D$. Intuitively, $M^{l_{i}}$ indicates the embedding landscape of the dataset on the $i$-th layer and the shape of the matrix is solely dependent on the size of the dataset. Similar embedding landscapes (i.e., similar matrix) in two layers indicate they extract comparable semantic information from inputs. With the matrix $M^{l}$ for each layer $l$, DeepArc then follows the centered kernel alignment (Kornblith et al., 2019) (CKA) metric to calculate the similarity between $M^{l_{i}}$ and $M^{l_{j}}$ to represent the semantic similarity between layer $f^{l_{i}}$ and $f^{l_{j}}$.
@@ 2.3. DNNs Slicing
DNNs, essentially programs made up of artificial neurons, have seen recent innovations in “slicing” techniques. These methods identify key neurons within DNNs that are crucial for various tasks (Xie et al., 2022; Zhang et al., 2020a). The essence of DNN slicing research revolves around using a neuron contribution metric to evaluate the impact of each neuron on a dataset $D$. A neuron selection strategy is then applied to pick out critical neurons based on their contribution to the network’s functionality.
Specifically, the process could be formulated as follows: we calculate $\mathbf{C}_{n}^{l,D}$ for each neuron $n$ in layer $l$ based on a dataset $D$, the neuron contribution of the layer $l$ composed of neuron set $\mathbf{N}^{l}$ could be represented as $\mathbf{C}^{l,D}=\{\mathbf{C}_{n_{0}}^{l,D},\mathbf{C}_{n_{1}}^{l,D},...,\mathbf{C}_{n_{|\mathbf{N}^{l}|-1}}^{l,D}\}$. We further determine the selection of critical neurons according to the calculated neuron contribution and denote the selected critical neurons as $\mathbf{\Omega}^{l,D}_{k}$:
(3) $\mathbf{\Omega}_{k}^{l,D}=\text{top}\-k(\mathbf{N}^{l,D},\mathbf{C}^{l,D})$
@@
The predominant approaches of neuron contribution metric design for neuron $n$ are as follows: ❶ Avg (Average of Neuron Activation (Pei et al., 2017)). We compute the mean activation value for each neuron across the dataset, defined as: $avg_{n}(D)=\frac{\sum_{x\in D}f_{n}(x)}{|D|}.$Var (Variance of Neuron Activation (Ish-Horowicz et al., 2019)). We compute the activation value variance for each neuron using the formula: $var_{n}(D)=\frac{\sum_{x\in D}(f_{n}(x)-\overline{f_{n}(x)})^{2}}{|D|}.$Ens (Ensemble of Neuron Activation). With $avg_{n}(D)$ and variance $var_{n}(D)$ of each neuron in the neural network concerning dataset $D$, we normalize the two values to [0,1]. ❹ DeepLIFT (Interpretability-based Score (Xie et al., 2022; Li et al., 2023)). Given its interpretability, our analysis predominantly utilizes DeepLIFT (Shrikumar et al., 2017) for computations. In specific, DeepLIFT assigns contribution scores to each neuron based on the gradient during the backward propagation process. ❺ Taylor (Pruning-oriented Score (Molchanov et al., 2019)). Taylor utilizes model weights and gradient information to assess the contribution of neurons.
Leveraging the current neuron contribution metrics, there are intuitively two neuron selection strategies. The global neuron selection method calculates neuron contribution scores across the entire dataset $D$, followed by a selection of a specific range of neurons. In contrast, the category-wise neuron selection approach computes scores for each category of data ($D_{c_{0}}$, $D_{c_{1}}$, $\ldots$) separately, identifies a range of critical neurons for each category and layer (i.e., identifying $\mathbf{\Omega}_{k}^{l,D_{c_{0}}},\mathbf{\Omega}_{k}^{l,D_{c_{1}}},\ldots$), and ultimately merges these category-wise critical neurons for the entire dataset as a whole ($\mathbf{\Omega}_{k}^{l,D}$).
Figure 1. Overview of NeuSemSlice Framework. In semantic slicing, neuron contributions are highlighted in $purple$, with darker shades signifying greater contributions. Critical neurons are indicated in $orange$. For semantic-aware model maintenance, we preserve only these critical neurons. During model re-adaptation, we train solely the critical neurons, marked in $green$. For incremental development, only the non-critical neurons are trained, shown in $blue$.

### v2

@@ Abstract.
Deep Neural networks (DNNs), extensively applied across diverse disciplines, are characterized by their integrated and monolithic architectures, setting them apart from conventional software systems. This architectural difference introduces particular challenges to maintenance tasks, such as model restructure (e.g., model compression), re-adaptation (e.g., fitting new samples), and incremental development (e.g., continual knowledge accumulation).
Prior research addresses these challenges by identifying task-critical neuron layers, and dividing neural networks into semantically-similar sequential modules. However, such layer-level approaches fail to precisely identify and manipulate neuron-level semantic components, restricting their applicability to finer-grained model maintenance tasks.
@@
For semantic-aware model maintenance tasks, we provide a series of novel strategies based on semantic slicing to enhance NeuSemSlice. They include semantic components (i.e., critical neurons) preservation for model restructure, critical neuron tuning for model re-adaptation, and non-critical neuron training for model incremental development. A thorough evaluation has demonstrated that NeuSemSlice significantly outperforms baselines in all three tasks.
Deep Neural Networks, Model Maintenance
copyright: acmlicensedjournal: TOSEMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/3731556ccs: Software and its engineering Maintaining softwareccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
In the evolving landscape of artificial intelligence, DNNs represent a dynamic shift towards data-driven programming paradigms, extensively adopted across various fields (Kollias et al., 2018; Bansal et al., 2019; Théate and Ernst, 2021). As these models increasingly serve as core components in modern software, their maintenance transcends mere optimization challenges and requires strategies grounded in established software engineering principles (Wei et al., 2023; Latendresse et al., 2024; Ren et al., 2023). Analogous to how traditional software evolves to meet shifting requirements and enhance functionality, effective maintenance of DNNs is crucial for combating model drift and ensuring sustained accuracy and reliability as they are continuously exposed to new and changing data.
Unlike traditional software systems characterized by their modular and discrete structures, DNNs are typically designed as cohesive, monolithic entities. The unique architecture introduces distinct and challenging maintenance tasks, including model restructure, re-adaptation, and incremental development. They are pivotal elements for the operational robustness and longevity of neural network-based systems in a rapidly advancing technological milieu (Chen and Das, 2023). By framing DNN maintenance as an extension of traditional software maintenance practices, we address a critical need within the software engineering community, ensuring that DNNs continue to function correctly, evolve smoothly, and remain aligned with evolving requirements.
@@
Despite various proposals in the literature for maintaining DNN models like the DeepArc framework (Ren et al., 2023), these methods continue to face several limitations. One major challenge with traditional strategies, which depend largely on periodic model retraining, is their ineffectiveness in combating the unavoidable problem of model drift. DeepArc addresses this problem by modularizing neural networks, organizing layers into modules based on their semantic similarities. This strategy streamlines maintenance tasks, facilitating the pruning and fine-tuning of specific modules or layers for more efficient updates. However, this approach is constrained by its focus on layer-level manipulation. A more refined strategy that decomposes semantic components at a finer level, such as individual neurons, could enhance model restructuring and adaptation. Specifically, selectively pruning and tuning a small number of neurons might offer better effectiveness. Furthermore, the incremental development of models (Wang et al., 2024b) requires preserving the semantics and functions of previous versions while introducing new features. Given that semantics and functions in DNNs are primarily structured around neurons (Du et al., 2019; Xie et al., 2022; Zhang et al., 2020b; Li et al., 2021; Shrikumar et al., 2017; Li et al., 2024b, 2023a, a), the method of identifying and updating whole layers falls short of preserving the original semantics or functions during the incremental development process.
@@
Acknowledging the existing limitations, our goal is to develop techniques that enhance DNN model maintenance. Drawing inspiration from the significant benefits of dynamic slicing in the maintenance and evolution of traditional software (Beszédes et al., 2001; Jiang et al., 2014), we adopt similar slicing techniques for DNNs (Zhang et al., 2020a; Li et al., 2023b; Xie et al., 2022; Sun et al., 2022; Liu et al., 2018). Specifically, dynamic slicing excels in identifying and isolating essential code segments for implementing changes without causing unintended effects. By analogy, DNN slicing provides a systematic way to identify and separate semantic components at the neuron level, mirroring how code slicing pinpoints essential segments that influence a program’s behavior. Focusing interventions on the neurons most critical to the model’s functionality ensures targeted changes without introducing unintended side effects in non-critical parts.
@@
The semantic slicing module is designed to accurately identify critical neurons across the entire search space by employing a category-specific [^1] To be precise, we identify critical neurons within the network model that have a significant impact on the output of each output neuron, where different categories refer to different output neurons. For example, in classification tasks, categories can be determined based on the labels in the dataset; in regression tasks, it can be considered a single-category task; in other types of tasks, different output neurons are treated as different categories. strategy for optimal performance. Specifically, we compute neuron contribution scores (Pei et al., 2017; Ish-Horowicz et al., 2019; Xie et al., 2022; Li et al., 2023b; Molchanov et al., 2019) for each neuron category individually, selecting a specific range of critical neurons for each category before combining them across the entire dataset. Given the potential variability in the optimal range of critical neurons across different categories and layers—which may hinder effective combination—we refine this process to focus on neurons that encapsulate the layer’s full semantic range. This is achieved by assessing the similarity between the characteristics of these chosen neurons and the entire layer. The subsequent merging strategy further enhances the utility and relevance of the identified critical neurons.
Mirroring and inspired by the traditional software engineering tasks, the semantic-aware model maintenance module utilizes semantic slicing, and provides the following maintenance tasks for DNN models:
Model Restructure: In traditional software maintenance, code slicing can be used to identify redundant or non-essential code, streamlining the program and improving efficiency (Komondoor and Horwitz, 2001). Similarly, in DNN model maintenance, model restructure focuses on retaining only the critical neurons necessary for maintaining predictive accuracy, while minimizing storage requirements. This step isolates non-essential components of the model, simplifying the overall structure by focusing on the critical parts.
Model Re-adaptation: Code slicing in software engineering helps identify and modify the core parts of a program when fixing bugs (Sintaha et al., 2023; Li et al., 2024c; Papotti et al., 2024). In DNNs, model re-adaptation focuses on retraining critical neurons to address prediction errors, reducing the need to adjust many parameters. This mirrors how targeted changes in code slicing resolve issues without overhauling the entire program.
Model Incremental Development: In traditional software maintenance, slicing can be used to store the historical semantic information of code, thereby assisting software evolution (Servant and Jones, 2012; Li, 2017). Similarly, model incremental development stabilizes critical neuron parameters to prevent catastrophic forgetting when integrating new knowledge. This ensures that core logic remains intact while new functionality is added seamlessly, akin to how code slicing maintains program stability during incremental updates.
@@
We conduct extensive experiments to evaluate the performance of NeuSemSlice. For model restructure, NeuSemSlice enhances the efficiency of existing model compression techniques by an average of 13.46%, while also improving compression rate. For model re-adaptation, NeuSemSlice only trains an average of 61.49% of the parameters, enhancing both accuracy and efficiency. For incremental model development, NeuSemSlice performs well, boosting average accuracy by approximately 40% compared to direct retraining.
@@
Contributions. The contributions of this paper is summarized as follows: ❶ We propose NeuSemSlice, including semantic slicing for finer-grained neuron-level decomposition, and semantic-aware model maintenance, offering novel and effective strategies for semantic-aware DNN model maintenance tasks. ❷ We detail a sophisticated semantic slicing technique that employs a category-specific strategy to accurately identify and combine critical neurons across the dataset, refining the process of neuron selection based on semantic range to optimize model performance. ❸ We tackle the challenging maintenance tasks via the semantic-aware model maintenance module, which significantly advances the tasks of model restructure, re-adaptation, and incremental development. ❹ We conduct extensive experiments on three model maintenance tasks, showcasing the superior performance of NeuSemSlice.
@@ 2.1. Deep Neural Networks
In this work, we mainly focus on the classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{X}$ is a set of inputs and $\mathcal{Y}$ is a set of classes. Given an input $x\in\mathcal{X}$, we use $f^{l}(x)$ to represent the internal features extracted by the layer $l$ and $f_{n}(x)$ to represent the internal features extracted by the neuron $n$ (i.e., the output values of neurons at $l$). $\hat{y}$ represents the output of DNN (i.e., $f(x)$) and $y$ is the label of the given input $x$.
@@ 2.2. DNN Model Maintenance
In the realm of software engineering, maintaining and evolving systems has long been recognized as a critical challenge, and DNNs are no exception. As these models increasingly integrate into complex, safety-critical, and rapidly changing environments, practitioners must apply well-established maintenance principles—such as iterative improvement, modular refactoring, and continuous quality assurance—to keep DNNs robust, adaptable, and high-performing. This concept, akin to the established practices of traditional software maintenance, addresses the evolving needs within the software engineering domain for more efficient and adaptable DNN systems.
DeepArc (Ren et al., 2023) has introduced a new neural network architecture strategy to cut model maintenance costs. This approach identifies semantically similar layers as redundant, allowing for model restructuring by pruning these layers. Only non-redundant layers are fine-tuned on the target data, based on calculated semantic similarities of layer representations: Given a dataset $D$, the matrix $M^{l_{i}}$ is denoted as the similarity matrix for the $i$-th network layer $l_{i}$, where
@@
with $x_{a}$ and $x_{b}$ representing any two inputs from the dataset $D$. Intuitively, $M^{l_{i}}$ indicates the embedding landscape of the dataset on the $i$-th layer and the shape of the matrix is solely dependent on the size of the dataset. Similar embedding landscapes (i.e., similar matrices) in two layers indicate they extract comparable semantic information from inputs. With the matrix $M^{l}$ for each layer $l$, DeepArc then follows the centered kernel alignment (Kornblith et al., 2019) (CKA) metric to calculate the similarity between $M^{l_{i}}$ and $M^{l_{j}}$ to represent the semantic similarity between layer $f^{l_{i}}$ and $f^{l_{j}}$.
@@ 2.3. DNNs Slicing
DNNs, essentially programs made up of artificial neurons, have seen recent innovations in “slicing” techniques. These techniques, although termed as “slicing,” essentially represent advanced methods for neuron selection. The key idea behind DNN slicing is to evaluate and identify the most influential neurons in a network for various tasks (Xie et al., 2022; Zhang et al., 2020a). In this context, DNN slicing is closely related to neuron contribution metrics, which quantify the impact of each individual neuron on the overall performance of the model, particularly on a dataset $D$. The essence of DNN slicing research is not only to partition the network but to systematically select neurons that significantly contribute to the network’s functionality, based on their contribution scores. A neuron selection strategy is then applied to identify and extract these critical neurons, thereby enabling focused modifications, optimizations, or analysis of the network’s performance.
Specifically, the process could be formulated as follows: we calculate $\mathbf{C}_{n}^{l,D}$ for each neuron $n$ in layer $l$ based on a dataset $D$. The neuron contribution of the layer $l$ composed of neuron set $\mathbf{N}^{l}$ could be represented as $\mathbf{C}^{l,D}=\{\mathbf{C}_{n_{0}}^{l,D},\mathbf{C}_{n_{1}}^{l,D},...,\mathbf{C}_{n_{|\mathbf{N}^{l}|-1}}^{l,D}\}$. We further determine the selection of critical neurons according to the calculated neuron contribution and denote the selected critical neurons as $\mathbf{\Omega}^{l,D}_{k}$:
(3) $\mathbf{\Omega}_{k}^{l,D}=\text{top}\-k(\mathbf{N}^{l},\mathbf{C}^{l,D})$
@@
The predominant approaches of neuron contribution metric design for neuron $n$ are as follows: ❶ Avg (Average of Neuron Activation (Pei et al., 2017)). We compute the mean activation value for each neuron across the dataset, defined as: $avg_{n}(D)=\frac{\sum_{x\in D}f_{n}(x)}{|D|}.$Var (Variance of Neuron Activation (Ish-Horowicz et al., 2019)). We compute the activation value variance for each neuron using the formula: $var_{n}(D)=\frac{\sum_{x\in D}(f_{n}(x)-\overline{f_{n}(x)})^{2}}{|D|}.$Ens (Ensemble of Neuron Activation). For each neuron, we compute its $avg_{n}(D)$ and $var_{n}(D)$ on dataset $D$, normalize both values to [0,1], and then combine them into a single ensemble-based contribution score. ❹ DeepLIFT (Interpretability-based Score (Xie et al., 2022; Li et al., 2023b)). Given its interpretability, our analysis predominantly utilizes DeepLIFT (Shrikumar et al., 2017) for computations. Specifically, DeepLIFT assigns contribution scores to each neuron based on the gradient during the backward propagation process. ❺ Taylor (Pruning-oriented Score (Molchanov et al., 2019)). Taylor utilizes model weights and gradient information to assess the contribution of neurons.
Leveraging the current neuron contribution metrics, there are intuitively two neuron selection strategies. The global neuron selection method calculates neuron contribution scores across the entire dataset $D$, followed by a selection of a specific range of neurons. In contrast, the category-wise neuron selection approach computes scores for each category of data ($D_{c_{0}}$, $D_{c_{1}}$, $\ldots$) separately, identifies a range of critical neurons for each category (i.e., identifying $\mathbf{\Omega}_{k}^{l,D_{c_{0}}},\mathbf{\Omega}_{k}^{l,D_{c_{1}}},\ldots$), and ultimately merges these category-wise critical neurons for the entire dataset as a whole ($\mathbf{\Omega}_{k}^{l,D}$).
Figure 1. Overview of NeuSemSlice Framework. In Semantic Slicing, Neuron Contributions Are Highlighted in $purple$, With Darker Shades Signifying Greater Contributions. Critical Neurons Are Indicated in $orange$. For Model Restructure, We Preserve Only These Critical Neurons. During Model Re-Adaptation, We Train Solely the Critical Neurons, Marked in $green$. For Model Incremental Development, Only the Non-Critical Neurons Are Trained, Shown in $blue$.